# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ 4.2 Label-Driven Admission and Forced Scheduling for Classification
Data: $(\ell_{k},\mu_{k},\sigma_{k}\}_{k\in\mathcal{K}},H$
1 $\gamma=\beta\leftarrow 1/\sqrt{K\ell_{\max}}$ // Label-driven & Optimistic Admission Parameters
@@ 3       Observe a new post $j(t)=j$ of type $k(j)=k$; if no new post, set $j(t)=\perp$
5       $\bar{r}_{k}(t)=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)$ where $\underline{h}_{k}(t),\bar{h}_{k}(t)$ are given by (11)
@@
Our main result (Theorem 3) is that, setting $\beta=\gamma=1/\sqrt{K\ell_{\max}}$, OLBACID achieves an average regret of $\tilde{O}(\sqrt{\ell_{\max}}+\ell_{\max}/\sqrt{T}+\ell_{\max}^{1.5}/T)$ when there is a large margin $\eta$ and small number of types $K$. In particular, the guarantee matches the lower bound $O(\sqrt{\ell_{\max}})$ when $T\geq\ell_{\max}$. Even if there is no margin (so classification is difficult), OLBACID still obtains an average regret of $\tilde{O}(\sqrt{\ell_{\max}}+\ell_{\max}/\sqrt{T}+\ell_{max}^{2}/T)$, which matches the lower bound when $T\geq\ell_{\max}^{1.5}$.
@@
To bound those three losses, we define the maximum per-period views and cost standard deviation as $v_{\max}=\max(1,\max_{k}v_{k})$, $\sigma_{\max}=\max\left(1,\max_{k\in\mathcal{K}}\sigma_{k}\right)$ respectively [^6] We take maximum with $1$ to simplify the exposition but this does not impose an assumption on the instance., and the minimum per-period review rate as $\hat{\mu}_{\min}=\min_{t\in[T],k\in\mathcal{K}}N(t)\mu_{k}$. To account for uncertainty, we take $r_{\max}=v_{\max}(H+\sigma_{\max})\geq 1$; this is an upper bound on $r_{k}^{O}(\cdot),r_{k}^{R}(\cdot)$ by their definition Eq. (5). We focus on $T\geq 3$ and $K\leq\ell_{\max}$ (the bound in Theorem 3 becomes trivial if $K>\ell_{\max}$). Our bounds (Lemmas 4.14.2 and 4.3) hold for general $\beta,\gamma$, and are proven in Sections 4.34.4 and 4.5 respectively. The lemmas are defined for $T\geq 3$ and $K\leq\ell_{\max}$ (the bound in Theorem 3 becomes trivial otherwise). The proof of Theorem 3 (provided in Appendix C.4) directly combines the lemmas.
@@
$2K\beta r^{2}_{\max}\ell_{\max}T+\frac{33r_{\max}K\sigma^{2}_{\max}\ln T}{\hat{\mu}_{\min}\max(\eta,\gamma)^{2}}.$
@@
$4\ell_{\max}\left(v_{\max}\gamma T+Kr_{\max}\right)+\frac{33r_{\max}\ell_{\max}K\sigma_{\max}^{2}\ln T}{\hat{\mu}_{\min}\max(\eta,\gamma)^{2}}.$
@@
$\mathcal{L}^{\star}(w,T)+\frac{wT}{\beta}+\frac{66r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}+18Kr_{\max}\ell_{\max}\left(\sqrt{T\ln T}+\beta r_{\max}\ell_{\max}\right)$
@@
Our second lemma applies concentration bounds (Appendix C.5) to show that the event $\mathcal{E}_{k,t}=\{h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]\}$ (expected cost lies in confidence interval) holds with high probability.
@@
For any $k,t$, the expected cost lies in $[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ with probability $\mathbb{P}\{\mathcal{E}_{k,t}\}\geq 1-2t^{-3}$.
@@
The label-driven queue admits at most $\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\right]\leq\frac{33K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}$ posts.
@@
$r_{\max}\mathbb{E}\left[\sum_{t=1}^{T}\left(\sum_{k\in\mathcal{K}}Q_{k}(t)+|\mathcal{Q}^{\mathcal{\textsc{LD}}}(t)|\right)\right]\leq 2r_{\max}K\beta r_{\max}\ell_{\max}T+\frac{33r_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}},$
@@
Our first lemma (proof in Appendix C.7) shows a Lipschitz property for functions $r_{k}^{O}(\cdot),r_{k}^{R}(\cdot),r_{k}(\cdot)$.
@@
For any type $k$, functions $r_{k}^{O}(h),r_{k}^{R}(h)$ and $r_{k}(h)$ are $1.5v_{k}$-Lipschitz continuous.
Our next lemma (proof in Appendix C.8) uses this property to establish a bound on the type-$k$ period-$t$ classification loss $Z_{k}(t)=(Y_{k}(t)(r_{k}^{O}-r_{k}^{R})^{+}+(1-Y_{k}(t))(r_{k}^{R}-r_{k}^{O})^{+})\ell_{k}(\Lambda_{k}(t)-A_{k}(t)-E_{k}(t))$.
Lemma 4.9.
For any type $k$ and period $t$, $Z_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\leq\left(3v_{\max}\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\Lambda_{k}(t).$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}Z_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+Kr_{\max}\ell_{\max}\sum_{t=1}^{T}\frac{2}{t^{3}}$ (By Lemma 4.5)
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\left(3v_{\max}\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\right]+4Kr_{\max}\ell_{\max}$ (By Lemma 4.9)
$\leq 3v_{\max}\gamma\ell_{\max}T+\frac{33r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}+4Kr_{\max}\ell_{\max}.$ (By Lemmas 4.44.6)
@@
The regret in admission is $\textsc{RegA}(T)\leq 12K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+6K\beta r_{\max}\ell_{\max}\sqrt{8T\ln T}.$
@@
The regret in scheduling is $\textsc{RegS}(T)\leq\frac{66\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$.
@@
$\leq wT+\beta\mathcal{L}^{\star}(w,T)+12K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+6K\beta r_{\max}\ell_{\max}\sqrt{8T\ln T}+\frac{66\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$ (By Lemmas 4.11 and 4.12)
$\leq wT+\beta\mathcal{L}^{\star}(w,T)+\frac{66\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}+18K\beta r_{\max}\ell_{\max}\left(\sqrt{T\ln T}+\beta r_{\max}\ell_{\max}\right).$
@@ 5 OLBACID with Type Aggregation and Contextual Learning
In this section, we design an algorithm whose performance does not deteriorate with the number of types $K$. We adopt a linear contextual structure assumption that is common in the bandit literature [LCLS10, CLRS11, APS11], and in online content moderation practice [ABB+22]. Each post comes with a $d$-dimensional feature vector $\boldsymbol{\phi}_{k}\in\mathbb{R}^{d}$ associated with its type $k$ and its expected cost satisfies a linear model $h_{k}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star}$ for a fixed unknown vector $\boldsymbol{\theta}^{\star}$.
@@
We incorporate the contextual information in our confidence intervals around $h_{k}$ by classical bandit techniques [APS11]; we provide a more detailed comparison in Appendix D.1. We construct a confidence set $\mathcal{C}_{t}$ for $\boldsymbol{\theta}^{\star}$ based on the collected data set $\mathcal{D}(t)$. Recalling that $\mathcal{S}(\tau)$ is the reviewed post for period $\tau$ (or $\perp$ when no post is reviewed), the dataset in period $t$ is $\{(\boldsymbol{X}_{\tau},Z_{\tau})\}_{\tau\leq t}$ with $\boldsymbol{X}_{\tau}=\boldsymbol{\phi}_{k(\mathcal{S}(\tau))}$ and $Z_{\tau}=c_{\mathcal{S}(\tau)}$. We use the ridge estimator with regularization parameter $\kappa$, i.e,
$\hat{\boldsymbol{\theta}}(t)=\bar{\boldsymbol{V}}_{t}^{-1}[\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{t}][Z_{1},\ldots,Z_{t}]^{\intercal}\quad\text{where}\quad\bar{\boldsymbol{V}}_{t}=\kappa\boldsymbol{I}+\sum_{\tau\leq t}\boldsymbol{X}_{\tau}\boldsymbol{X}_{\tau}^{\intercal}\quad\text{and}\quad\hat{\boldsymbol{\theta}}(0)=\boldsymbol{0}.$ (13)
@@
Recalling that $\sigma_{\max}=\max_{k\in\mathcal{K}}\sigma_{k}$ is the maximum idiosyncrasy variance and $U$ upper bounds the Euclidean norm of $\boldsymbol{\theta}^{\star}$ and any feature $\boldsymbol{\phi}_{k}$, we define the confidence set with a confidence level $\delta$ by
@@
$\mathcal{C}_{t}\coloneqq\left\{\boldsymbol{\theta}\in\mathbb{R}^{d}\colon\|\hat{\boldsymbol{\theta}}(t)-\boldsymbol{\theta}\|_{\bar{\boldsymbol{V}}_{t}}\leq B_{\delta}(t)\right\}\quad\text{where }B_{\delta}(t)\coloneqq\sigma_{\max}\sqrt{d\ln\left(\frac{1+tU^{2}/\kappa}{\delta}\right)}+\sqrt{\kappa}U.$ (14)
Using confidence set $\mathcal{C}_{t-1}$, we can modify our confidence intervals from (11) for period $t$ as follows:
$\underline{h}_{k}(t)=\max\left(-H,\min_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}\right),\quad\bar{h}_{k}(t)=\min\left(H,\max_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}\right),\quad\bar{r}_{k}(t)=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h).$ (15)

### v3

@@ 4.2 Label-Driven Admission and Forced Scheduling for Classification
Data: $(\ell_{k},\mu_{k}\}_{k\in\mathcal{K}},\sigma_{\max},r_{\max}$     Admission parameters: $\gamma=\beta\leftarrow 1/\sqrt{K\ell_{\max}}$
@@ 2       Observe a new post $j(t)=j$ of type $k(j)=k$; if no new post, set $j(t)=\perp$
4       Calculate $\underline{h}_{k}(t),\bar{h}_{k}(t),\bar{r}_{k}(t)$ by (11) and (12)
@@
Our main result (Theorem 3) is that, setting $\beta=\gamma=1/\sqrt{K\ell_{\max}}$, OLBACID achieves an average regret of $\tilde{O}(\sqrt{\ell_{\max}}+\ell_{\max}/\sqrt{T}+\ell_{\max}^{1.5}/T)$ when there is a large margin $\eta\coloneqq\min(1,\min_{k\in\mathcal{K}}|h_{k}|)$ and small number of types $K$. In particular, the guarantee matches the lower bound $O(\sqrt{\ell_{\max}})$ when $T\geq\ell_{\max}$. Even if there is no margin (so classification is difficult), OLBACID still obtains an average regret of $\tilde{O}(\sqrt{\ell_{\max}}+\ell_{\max}/\sqrt{T}+\ell_{max}^{2}/T)$, which matches the lower bound when $T\geq\ell_{\max}^{1.5}$.
@@
To bound the losses, we define the minimum per-period review rate as $\hat{\mu}_{\min}=\min_{t\in[T],k\in\mathcal{K}}N(t)\mu_{k}$. We focus on $T\geq 3$ and $K\leq\ell_{\max}$ (the bound in Theorem 3 becomes trivial if $K>\ell_{\max}$). Our bounds (Lemmas 4.14.2 and 4.3) hold for general $\beta,\gamma$, and are proven in Sections 4.34.4 and 4.5 respectively. The lemmas are defined for $T\geq 3$ and $K\leq\ell_{\max}$ (the bound in Theorem 3 becomes trivial otherwise). The proof of Theorem  3 (provided in Appendix C.3) directly combines the lemmas.
@@
$2K\beta r^{2}_{\max}\ell_{\max}T+\frac{38r_{\max}K\sigma^{2}_{\max}\ln T}{\hat{\mu}_{\min}\max(\eta,\gamma)^{2}}.$
@@
$5\ell_{\max}\left(\gamma T+Kr_{\max}\right)+\frac{38r_{\max}\ell_{\max}K\sigma_{\max}^{2}\ln T}{\hat{\mu}_{\min}\max(\eta,\gamma)^{2}}.$
@@
$\mathcal{L}^{\star}(w,T)+\frac{wT}{\beta}+\frac{76r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}+20Kr_{\max}\ell_{\max}\left(\sqrt{T\ln T}+\beta r_{\max}\ell_{\max}\right)$
@@
Our second lemma applies concentration bounds (Appendix C.4) to show that the event $\mathcal{E}_{k,t}=\{h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)],r_{k}\leq\bar{r}_{k}(t)\}$ (confidence bounds are valid) holds with high probability. A challenge in establishing this lemma is to show that $c_{j}^{+}$ and $c_{j}^{-}$ are also sub-Gaussian given that $c_{j}$ is sub-Gaussian, which we establish with techniques from [Kon14].
@@
For any $k,t$, the confidence bounds in (11),(12) are valid with probability $\mathbb{P}\{\mathcal{E}_{k,t}\}\geq 1-4t^{-3}$.
@@
The label-driven queue admits at most $\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\right]\leq\frac{38K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}$ posts.
@@
$r_{\max}\mathbb{E}\left[\sum_{t=1}^{T}\left(\sum_{k\in\mathcal{K}}Q_{k}(t)+|\mathcal{Q}^{\mathcal{\textsc{LD}}}(t)|\right)\right]\leq 2r_{\max}K\beta r_{\max}\ell_{\max}T+\frac{38r_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}},$
@@
The lemma below (proof in Appendix C.6) establishes a bound on the type-$k$ period-$t$ classification loss $Z_{k}(t)=(Y_{k}(t)(r_{k}^{O}-r_{k}^{R})^{+}+(1-Y_{k}(t))(r_{k}^{R}-r_{k}^{O})^{+})\ell_{k}(\Lambda_{k}(t)-A_{k}(t)-E_{k}(t))$.
@@
For any type $k$ and period $t$, $Z_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\leq\left(\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\Lambda_{k}(t).$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}Z_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+Kr_{\max}\ell_{\max}\sum_{t=1}^{T}\frac{4}{t^{3}}$ (By Lemma 4.5)
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\left(\gamma\ell_{\max}+r_{\max}\ell_{\max}Q^{\mathcal{\textsc{LD}}}(t)\right)\right]+5Kr_{\max}\ell_{\max}$ (By Lemma 4.8)
$\leq\gamma\ell_{\max}T+\frac{38r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}+5Kr_{\max}\ell_{\max}.$ (By Lemmas 4.44.6)
@@
The regret in admission is $\textsc{RegA}(T)\leq 20K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+8K\beta r_{\max}\ell_{\max}\sqrt{8T\ln T}.$
@@
The regret in scheduling is $\textsc{RegS}(T)\leq\frac{76\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$.
@@
$\leq wT+\beta\mathcal{L}^{\star}(w,T)+20K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+8K\beta r_{\max}\ell_{\max}\sqrt{T\ln T}+\frac{76\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$ (By Lemmas 4.10 and 4.11)
$\leq wT+\beta\mathcal{L}^{\star}(w,T)+\frac{76\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}+20K\beta r_{\max}\ell_{\max}\left(\sqrt{T\ln T}+\beta r_{\max}\ell_{\max}\right).$
@@ 5 OLBACID with Type Aggregation and Contextual Learning
In this section, we design an algorithm whose performance does not deteriorate with the number of types $K$. We adopt a linear contextual structure assumption that is common in the bandit literature [LCLS10, CLRS11, APS11], and in online content moderation practice [ABB+22]. In particular, we assume that each post comes with a $d$-dimensional feature vector $\boldsymbol{\phi}_{k}\in\mathbb{R}^{d}$ associated with its type $k$. The expected cost of accepting a post, $r_{k}^{O}=\mathbb{E}_{k}\left[c_{j}^{+}\right]$, and the expected cost of rejecting a post, $r_{k}^{R}=\mathbb{E}_{k}\left[-c_{j}^{-}\right]$, both satisfy a linear model such that $r_{k}^{O}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star,O}$ and $r_{k}^{R}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star,R}$ with two fixed unknown vectors $\boldsymbol{\theta}^{\star,O},\boldsymbol{\theta}^{\star,R}\in\mathbb{R}^{d}$. This model also implies $h_{k}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star,h}$ with $\boldsymbol{\theta}^{\star,h}=\boldsymbol{\theta}^{\star,O}-\boldsymbol{\theta}^{\star,R}$. Letting $\boldsymbol{R}_{k}$ be the vector $[r_{k}^{O},r_{k}^{R}]$ and $\boldsymbol{\Theta}^{\star}=[\boldsymbol{\theta}^{\star,O},\boldsymbol{\theta}^{\star,R}]$, the linear model is equivalent to $\boldsymbol{R}_{k}=\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\Theta}^{\star}.$ We assume that the Euclidean norms of $\boldsymbol{\theta}^{\star,O},\boldsymbol{\theta}^{\star,R}$ and feature vectors $\boldsymbol{\phi}_{k}$ are upper bounded by a known constant $U\geq 1$.
@@
We incorporate the contextual information in our confidence intervals around $h_{k}$ by classical bandit techniques [APS11]; we provide a more detailed comparison in Appendix D.1. We construct a confidence set $\mathcal{C}_{t}$ for $\boldsymbol{\Theta}^{\star}$ based on the collected data set $\mathcal{D}(t)$. Recalling that $\mathcal{S}(\tau)$ is the reviewed post for period $\tau$ (or $\perp$ when no post is reviewed), the dataset in period $t$ is $\{(\boldsymbol{X}_{\tau},\boldsymbol{Z}_{\tau})\}_{\tau\leq t}$ with $\boldsymbol{X}_{\tau}=\boldsymbol{\phi}_{k(\mathcal{S}(\tau))}$ and $\boldsymbol{Z}_{\tau}=\begin{pmatrix}c^{+}_{\mathcal{S}(\tau)}\\ c^{-}_{\mathcal{S}(\tau)}\end{pmatrix}$. We use the ridge estimator with regularization parameter $\kappa$,
$\hat{\boldsymbol{\Theta}}(t)=[\hat{\boldsymbol{\theta}}^{O}(t),\hat{\boldsymbol{\theta}}^{R}(t)]$ $=\bar{\boldsymbol{V}}_{t}^{-1}[\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{t}][\boldsymbol{Z}_{1},\ldots,\boldsymbol{Z}_{t}]^{\intercal}$ (14)
$\text{where}~{}\bar{\boldsymbol{V}}_{t}=\kappa\boldsymbol{I}+\sum_{\tau\leq t}\boldsymbol{X}_{\tau}\boldsymbol{X}_{\tau}^{\intercal},~{}\hat{\boldsymbol{\Theta}}(0)=[\boldsymbol{0},\boldsymbol{0}].$
We also construct an estimator for $\boldsymbol{\theta}^{\star,h}$ by $\hat{\boldsymbol{\theta}}^{h}(t)=\hat{\boldsymbol{\theta}}^{O}(t)-\hat{\boldsymbol{\theta}}^{R}(t).$
@@
Recalling that $\sigma_{\max}$ is the variance proxy of cost distributions and $U$ upper bounds the Euclidean norm of $\boldsymbol{\theta}^{\star,O},\boldsymbol{\theta}^{\star,R}$ and any feature $\boldsymbol{\phi}_{k}$, we define the confidence sets with a confidence level $\delta$ by
@@
$\mathcal{C}^{O}_{t}\coloneqq\left\{\boldsymbol{\theta}^{O}\in\mathbb{R}^{d}\colon\|\hat{\boldsymbol{\theta}}^{O}(t)-\boldsymbol{\theta}^{O}\|_{\bar{\boldsymbol{V}}_{t}}\leq B_{\delta}(t)\right\},$ $\quad\mathcal{C}^{R}_{t}\coloneqq\left\{\boldsymbol{\theta}^{R}\in\mathbb{R}^{d}\colon\|\hat{\boldsymbol{\theta}}^{R}(t)-\boldsymbol{\theta}^{R}\|_{\bar{\boldsymbol{V}}_{t}}\leq B_{\delta}(t)\right\}$ (15)
where $\text{where }B_{\delta}(t)$ $\coloneqq\sigma_{\max}\sqrt{2d\ln\left(\frac{1+tU^{2}/\kappa}{\delta}\right)}+\sqrt{\kappa}U.$
For ease of notation, we also define the aggregated confidence set $\mathcal{C}_{t}=\mathcal{C}^{O}_{t}\times\mathcal{C}^{R}_{t}$ and the confidence set of $\boldsymbol{\theta}^{\star}_{h}$: $\mathcal{C}^{h}_{t}=\{\boldsymbol{\theta}^{O}-\boldsymbol{\theta}^{R}\colon[\boldsymbol{\theta}^{O},\boldsymbol{\theta}^{R}]\in\mathcal{C}\}$. Using confidence sets from period $t-1$, we can modify our confidence bounds from (11) and (12) for period $t$: :
$\underline{h}_{k}(t)=\max\left(-r_{\max},\min_{\boldsymbol{\theta}^{h}\in\mathcal{C}^{h}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{h}\right),~{}\bar{h}_{k}(t)=\min\left(r_{\max},\max_{\boldsymbol{\theta}^{h}\in\mathcal{C}^{h}_{t-1}}\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{h}\right).$ (16)
$\bar{r}_{k}(t)=\max_{[\boldsymbol{\theta}^{O},\boldsymbol{\theta}^{R}]\in\mathcal{C}_{t-1}}\min\left\{\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{O},\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{R}\right\}.$ (17)